OpenStudy (anonymous):

Compute and write your answer in scientific notation. (4.3 x 10-8) (1.5 x 109)

OpenStudy (anonymous):

@BloomLocke367 can you hel p

OpenStudy (anonymous):

(4.3 x 10-8) (1.5 x 109) 4.3*1.5=6.45 10-8*109=10 4.3*1.5*10=64.5

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (anonymous):

@leader can you help

OpenStudy (anonymous):

@TheSmartOne can you help?

OpenStudy (bloomlocke367):

sorry... I had to leave to eat dinner... do you still need help?

OpenStudy (anonymous):

yes please I have this question and one more if you could help:]

OpenStudy (bloomlocke367):

okay.... you were on the right track.. but it asked for scientific notation... so you don't actually multiply the 10 out... make sense?

OpenStudy (anonymous):

I just put4.3*1.5*10

OpenStudy (bloomlocke367):

you multiply the 4.3 and 1.5... but not the 10

OpenStudy (anonymous):

6.45*10

OpenStudy (anonymous):

is that right?

OpenStudy (bloomlocke367):

yep :)

OpenStudy (anonymous):

cool thanks much! can you help me with one more?

OpenStudy (bloomlocke367):

\(\Huge\scr\color{lime}{Great~Job!}\) and of course cx

OpenStudy (anonymous):

8.2/2.75=2.98181 105/10-3=108 2.981*108

OpenStudy (bloomlocke367):

sorry, you need to open a new question.. I'll help you there

OpenStudy (anonymous):

k will do:]

OpenStudy (bloomlocke367):

okay. just tag me cx